### v1

# FPC-VLA: A Vision-Language-Action Framework with a Supervisor for Failure Prediction and Correction

Abstract
Robotic manipulation is a fundamental component of automation. However, traditional perception-planning pipelines often fall short in open-ended tasks due to limited flexibility, while the architecture of a single end-to-end Vision-Language-Action (VLA) offers promising capabilities but lacks crucial mechanisms for anticipating and recovering from failure. To address these challenges, we propose FPC-VLA, a dual-model framework that integrates VLA with a supervisor for failure prediction and correction. The supervisor evaluates action viability through vision–language queries and generates corrective strategies when risks arise, trained efficiently without manual labeling. A similarity-guided fusion module further refines actions by leveraging past predictions. Evaluation results on multiple simulation platforms (SIMPLER and LIBERO) and robot embodiments (WidowX, Google Robot, Franka) show that FPC-VLA outperforms state-of-the-art models in both zero-shot and fine-tuned settings. By activating the supervisor only at keyframes, our approach significantly increases task success rates with minimal impact on execution time. Successful real-world deployments on diverse, long-horizon tasks confirm FPC-VLA’s strong generalization and practical utility for building more reliable autonomous systems. Our project page is https://fpcvla.github.io/.
I Introduction
Robot manipulation is crucial for advancing automation in industries like manufacturing, logistics and healthcare [1, 2]. Traditional approaches follow a two-stage “perception-planning” pipeline: estimating object poses and predefined grasp points [3, 4], then generating motion trajectories. This requires explicit recognition and geometric modeling, assuming prior knowledge of target objects, which limits flexibility in open-ended tasks. Additionally, pose estimation and grasp-point labeling often fail under occlusion, clutter, or with novel objects, causing errors in perception that propagate to planning and lead to unreliable actions.
Unlike traditional modular pipelines, Vision-Language-Action (VLA) models [5, 6, 7, 8, 9, 10, 11, 12, 13] unify perception, task understanding, and motion generation into a single model, directly mapping visual observations and natural language instructions into low-level robot actions. By leveraging large-scale pretraining on vision-language data, such models can generalize across diverse objects, environments, and ambiguous commands without relying on explicit 3D modeling, which enables zero-shot or few-shot adaptation to new tasks. Additionally, the model dynamically infers tasks from language, enhancing intuitive human-robot interaction.
Figure 1: We propose FPC-VLA, a dual-model framework that integrates VLA with a supervisor for failure prediction and correction. It outperforms other advanced methods in all benchmarks.
Despite the significant progress in VLA models, current frameworks still face several notable limitations. First, as shown in Fig. 1(a), most existing approaches [5, 6, 7, 8, 9, 11, 12, 13, 10] rely heavily on imitation learning from expert demonstrations, which are typically composed only of successful trajectories. This leads to a fundamental weakness: these models struggle to recover from deviations, often requiring external supervision for potential failure prediction and corrective guidance. Moreover, collecting large-scale datasets that capture both failure prediction and correction is another bottleneck. Most existing datasets usually annotate failures post hoc [14, 15, 16, 17], and rarely enable early prediction or proactive interventions. The lack of a standardized or scalable methodology limits the model’s generalization ability. In addition, as shown in Fig. 1(b), many current methods segment actions into action chunks, losing the information of the latter part of the predicted sequence, and often overlook the multimodal nature of action trajectories—i.e., the fact that there may be multiple plausible ways to achieve a goal.
To address these challenges, we propose FPC-VLA, a vision-language-action framework with a supervisor for failure prediction and correction. A VLM-based supervisor verifies predicted actions before execution and, if necessary, provides directional corrections to refine motion. We also introduce an automated method to generate a large-scale failure prediction and correction dataset from RLDS-format data, pairing images, task instructions, and gripper-focused QA prompts. Additionally, a similarity-guided action fusion module is proposed to aggregate historical predictions using cosine similarity and temporal decay, producing smoother and more robust action outputs.
The main contributions of this work are as follows:
  1. 1.
    We propose a VLM-based supervisor that predicts and corrects potential robot failures by evaluating gripper actions using structured prompts and image inputs.
  2. 2.
    We introduce a similarity-guided action fusion module that aggregates historical predictions to produce smoother and more reliable robotic motions.
  3. 3.
    We evaluate FPC-VLA on SIMPLER and LIBERO benchmarks, using WidowX, Google Robot and Franka, and test it for real-world applicability on real robots. The results show that FPC-VLA outperforms other state-of-the-art methods.
II Related Work
II-A Vision-Language-Action Models
Recent advances in multi-modal learning have driven the development of VLAs. Notable early examples include RT-1 [5], RT-2 [6], and RT-X [7], which enable end-to-end training on large datasets, achieving strong generalization across platforms. Open-source models like Octo [8] and OpenVLA [9] have also emerged, pre-trained on extensive robot datasets and fine-tuned with domain-specific data. More recent efforts, such as SpatialVLA [10] with Ego3D encoding and adaptive action grids, and TraceVLA [11] with visual trace prompting, enhance spatial and spatiotemporal reasoning. Other developments include RoboVLMs [12], which use continuous outputs and policy heads, and CogACT [13], which employs diffusion action transformers for modeling action sequences. Despite these advances, VLA models still struggle with autonomously detecting and recovering from failures in unstructured environments.
II-B Robot Failure Detection and Correction
To address robot failure challenges, recent studies have explored using Multimodal Large Language Models (MLLMs) as auxiliary agents for error detection and reasoning. [14] can identify the reasons for failure, but does not provide a correction strategy. [18] implements failure correction but relies on human-provided instructions. [15, 16, 17] are able to autonomously adjust strategies after grasping failures, but lack the ability to anticipate such failures beforehand. Among them, RACER [16] also uses an external supervisor, but it only provides a corrected task instruction in natural language after a failure, which cannot be directly mapped to action changes. EvolvingGrasp [19] improves through preference alignment from both successes and failures, and Phoenix [20] connects semantic reflection with motion-level correction. Unlike the approaches primarily focusing on post-failure adaptations or manual intervention, our approach anticipates failure scenarios and provides preemptive corrective actions.
III Method
Figure 2: Architecture of FPC-VLA. The input to the framework consists of the observed image and a natural language instruction, and the output is the end-effector pose increment and the gripper state of the robotic arm.
III-A Architecture Overview and Problem Formulation
As shown in Fig. 2, the input to FPC-VLA consists of an RGB image $I_{t}\in\mathbb{R}^{H\times W\times 3}$ and a language instruction $l$ at timestep $t$ . The output is a refined action $\mathbf{a}^{{}^{\prime}}_{t}\in\mathbb{R}^{7}$ that corrects failure-prone predictions. The proposed pipeline includes three modules: 1) a robot failure dataset generation method, 2) a VLM-based supervisor for failure prediction and correction, and 3) a similarity-guided action fusion module.
In the dataset generation module, grasp events are identified by detecting changes in gripper state $g_{t}$, forming a change set $\mathcal{C}$. For each event $c\in\mathcal{C}$, we retain frames $\mathcal{R}_{c}$, and compute pose differences $\Delta\mathbf{a}$ . These are converted into structured language descriptions using thresholds $\delta_{\text{trans}}^{\text{small}},\delta_{\text{trans}}^{\text{large}},\delta_{\text{rot}}^{\text{small}},\delta_{\text{rot}}^{\text{large}}$, and combined with task and answer constraints to form prompts $\mathcal{P}_{t}$. In the similarity-guided action fusion module, a set of past predictions $\mathcal{A}_{t}$ is collected, each decomposed as $\mathbf{a}_{t}^{(k)}=[\mathbf{p}_{t}^{(k)},g_{t}^{(k)}]$ in observation $\mathcal{O}_{t-k}$. Cosine similarity $\text{sim}^{(k)}$ with the latest pose $\mathbf{p}_{t}^{(0)}$ is smoothed via sigmoid $s^{(k)}$, weighted by time decay $d^{(k)}$, and normalized to produce fusion weights $w^{(k)}$, yielding the fused action $\hat{\mathbf{a}}_{t}=[\hat{\mathbf{p}}_{t},\hat{g}_{t}]$. In the VLM-based failure prediction and correction module, when $g_{t}$ changes, the model $\mathcal{M}$ takes image $I_{t}$ and prompt $\mathcal{P}_{t}$ as input, returning response $R_{t}=\mathcal{M}(I_{t},\mathcal{P}_{t})$. If $R_{t}$ starts with “No”, translation correction $\Delta\mathbf{p}\in\mathbb{R}^{3}$ and rotation correction $\Delta r_{z}\in\mathbb{R}$ are extracted to obtain refined action $\mathbf{a}^{\prime}_{t}$.
III-B Vision-Language-Action Model
Vision Module  The vision module is responsible for extracting patch-level features from input images and projecting them into the embedding space of the language model. Given image $I_{t}$, it is passed through hybrid visual encoder built on DINOv2 [21] and SigLIP [22]. The encoded features are then passed through a linear projection to align with the LLM embedding space to get vision tokens $\mathcal{V}=\{v_{1},v_{2},...,v_{N_{\mathcal{V}}}\}$.
Language Module  The language module is built on Llama2 [23], a transformer-based architecture designed for autoregressive language modeling. The input consists of a sequence of language tokens $\mathcal{L}=\{l_{1},l_{2},\ldots,l_{N_{\mathcal{L}}}\}$, where each token $l_{i}$ is mapped to a high-dimensional embedding space $\mathbb{R}^{d}$. For multimodal inputs, the text embeddings $\mathcal{L}$ are combined with visual embeddings $\mathcal{V}$ from the vision module. The output cognition token $\mathcal{Z}\in\mathbb{R}^{d}$ is used as input for the action module to interpret and generate the required actions.
Action Module  The action module uses Diffusion Transformer (DiT) [24] to generate action sequences conditioned on $\mathcal{Z}$. Diffusion model generates actions by gradually denoising a noisy input, while the transformer backbone enhances the model’s ability to capture long-range dependencies in the data. During the training phase, the noisy sample $x_{t}$ is obtained using the q-sampling process
$x_{t}=\sqrt{\alpha_{t}}\cdot x_{0}+\sqrt{1-\alpha_{t}}\cdot\epsilon,$ (1)
where $\alpha_{t}\in(0,1)$ controls the noise level at timestep $t$, and $x_{0}$ represents the ground truth actions for the next $n$ steps. The model DiT is trained to predict the noise $\epsilon$ added to $x_{0}$, given the noisy input $x_{t}$, the timestep $t$, and the condition $\mathcal{Z}$
$\text{DiT}(x_{t},t,z)\approx\epsilon.$ (2)
III-C Robot Failure Prediction and Correction Dataset
Most existing large-scale VLA datasets [7, 25, 5, 26] only include successful trajectories, while current failure correction datasets [15, 17] are limited in scale and lack predictability. To address this, we propose an automated method for generating failure correction datasets from large-scale RLDS-format data and adapt them for VLM training.
Denote the dataset $\mathcal{D}$ as a collection of episodes. Each episode $\mathcal{E}_{i}$ consists of a sequence of steps
$\mathcal{D}=\{\mathcal{E}_{1},\mathcal{E}_{2},\dots,\mathcal{E}_{N}\},\ \mathcal{E}_{i}=\{s_{1},s_{2},\dots,s_{T}\},$ (3)
where each step $s_{t}$ includes:
  • An RGB image observation $I_{t}\in\mathbb{R}^{H\times W\times 3}$,
  • An action vector $\mathbf{a}_{t}\in\mathbb{R}^{7}$ (6-DOF pose + 1-D gripper),
  • A task-level language instruction $l$.
Discrete grasp events are detected by tracking changes in the gripper state $g_{t}$. The set of change timesteps is
$\mathcal{C}=\{t\mid|g_{t}-g_{t-1}|>\delta_{g}\},$ (4)
where the threshold $\delta_{g}$ is 0.5. For each change timestep $c\in\mathcal{C}$, retain a set of three preceding frames and itself
$\mathcal{R}_{c}=\{c-3,c-2,c-1,c\}\cap[1,T].$ (5)
The total set of retained indices is $\mathcal{R}=\bigcup_{c\in\mathcal{C}}\mathcal{R}_{c}$.
For every retained frame $s_{t}$ with $t\in\mathcal{R}$, we identify the next change timestep $c\in\mathcal{C}$ such that $c\geq t$, and compute the pose difference
$\Delta\mathbf{a}=\mathbf{a}_{c}-\mathbf{a}_{t}.$ (6)
For each $\Delta\mathbf{a}$, we construct a textual description based on movement and rotation direction. For each axis $i\in\{x,y,z\}$, a movement direction and magnitude are defined as
$T_{i}=\left\{\begin{array}[]{l r}\text{None},&|\Delta i|\leq\delta_{\text{trans}}^{\text{small}}\\ \text{Move $D$. Small},&\delta_{\text{trans}}^{\text{small}}<|\Delta i|\leq\delta_{\text{trans}}^{\text{large}}\\ \text{Move $D$. Large},&|\Delta i|>\delta_{\text{trans}}^{\text{large}}\end{array}\right.$.
Around the vertical axis $z$, a rotation direction is defined similarly as
$R=\left\{\begin{array}[]{l r}\text{None},&|\Delta r_{z}|\leq\delta_{\text{rot}}^{\text{small}}\\ \text{Rotate $D$. Small},&\delta_{\text{rot}}^{\text{small}}<|\Delta r_{z}|\leq\delta_{\text{rot}}^{\text{large}}\\ \text{Rotate $D$. Large},&|\Delta r_{z}|>\delta_{\text{rot}}^{\text{large}}\end{array}\right$.,
where $\delta_{\text{trans}}^{\text{small}},\delta_{\text{trans}}^{\text{large}},\delta_{\text{rot}}^{\text{small}},\delta_{\text{rot}}^{\text{large}}$ are thresholds that limit the magnitude of movement or rotation, $\Delta i=i_{t+1}-i_{t}$ represents the position change along the $i\text{-th}$ axis, and $D$ is the direction. Table I illustrates the correspondence between variables and directions using the BridgeV2 [25] dataset as an example.
TABLE I: Mapping of variables and directions in BridgeV2 dataset
Axis Variable Sign Direction $D$
X $\Delta x$ + forward
- backward
Y $\Delta y$ + left
- right
Z $\Delta z$ + up
- down
Around Z $\Delta r_{z}$ + counterclockwise
- clockwise
Given the retained frame $s_{t}$, a standard prompt $\mathcal{P}_{t}$ consists of three parts: task requirements, answer limitations, and example samples. An example is shown in Table II.
TABLE II: Example of failure prediction and correction prompts
Part Content
Task Requirements My task is to pick rxbar chocolate from bottom drawer and place on counter. Should the robot arm close its gripper in the next step? If not, please specify the direction and magnitude of the gripper’s movement and rotation.
Answer Limitations The direction must be chosen from the following options: move up, move down, move left, move right, rotate clockwise, and rotate counterclockwise. At most one option can be selected from each pair: move up or move down, move left or move right, and rotate clockwise or rotate counterclockwise. The magnitude should be either large or small. Your response should start with ‘Yes’ or ‘No’. If the answer is ‘No’, then include the direction and magnitude of both movement and rotation.
Example Samples Here are some example responses: Example 1: Yes. Example 2: No. Move up. Large. Move left. Small. Example 3: No. Move down. Large. Rotate clockwise. Small.
III-D A VLM-Based Supervisor for Robot Failure Prediction and Correction
Existing robotic manipulation systems based on VLA models often directly execute predicted commands without assessing their contextual appropriateness. These systems typically lack the ability to predict the consequences of actions or to autonomously correct erroneous decisions. To address this limitation, we propose a failure prediction and correction supervisor that evaluates the suitability of gripper actions before execution.
Let the predicted action at timestep $t$ be denoted as:
$\hat{\mathbf{a}}_{t}=[x_{t},y_{t},z_{t},r_{x,t},r_{y,t},r_{z,t},g_{t}],$ (9)
where $(x,y,z)\in\mathbb{R}^{3}$ is the desired end-effector position delta. $(r_{x},r_{y},r_{z})\in\mathbb{R}^{3}$ is the rotation delta. $g$ is the gripper command (0 for close, $1$ for open). Here, $\hat{\mathbf{a}}_{t}$ comes from the output of the action fusion module in Section III-E.
The framework checks gripper consistency by comparing the current $g_{t}$ and previous $g_{t-1}$. The robot failure prediction and correction supervisor is triggered when the difference exceeds the threshold $\delta_{g}$
$\left|g_{t}-g_{t-1}\right|>\delta_{g}.$ (10)
When the gripper state $g_{t}$ meets the condition in (10), the prompt $\mathcal{P}_{t}$ and the current image $I_{t}$ are input into a vision language model $\mathcal{M}$
$R_{t}=\mathcal{M}(I_{t},P_{t}),$ (11)
where $R_{t}$ represents the generated natural language response by the model $\mathcal{M}$.
If the response $R_{t}$ is “Yes.”, it indicates that $\mathcal{M}$ approves the decision made by VLA regarding the opening or closing of the gripper, and the action will continue. If $R_{t}$ begins with “No”, the action $\hat{\mathbf{a}}_{t}$ needs to be adjusted according to the modification suggestions provided by $\mathcal{M}$. In the latter case, we extract structured directives from the text, including translation correction $\Delta\mathbf{p}\in\mathbb{R}^{3}$ and rotation correction $\Delta r_{z}\in\mathbb{R}$, based on predefined directional mappings. Each directional token is paired with a discrete magnitude indicator (small or large), which is mapped to a scalar step size ($\delta_{\text{trans}}^{\text{small}}$ or $\delta_{\text{trans}}^{\text{large}}$, $\delta_{\text{rot}}^{\text{small}}$ or $\delta_{\text{rot}}^{\text{large}}$). Based on the language-guided correction, we generate a refined action
$\mathbf{a}^{\prime}_{t}=\hat{\mathbf{a}}_{t}+[\Delta x,\Delta y,\Delta z,0,0,\Delta r_{z},0],$ (12)
leaving unaffected components unchanged. The refined action $\mathbf{a}^{\prime}_{t}$ is then executed within the environment.
III-E Similarity-Guided Action Fusion Module
Figure 3: Simulation results of FPC-VLA on different robots. The model performs well even in long-horizon tasks and non-pick-and-place tasks (e.g., opening drawers and turning on the stove).
To obtain a robust and smooth action output and avoid waste in action prediction, we employ an action fusion strategy. At each timestep $t$, the output action $\hat{\mathbf{a}}_{t}$ is not solely determined by the most recent policy output, but rather by fusing a set of historical predictions, each made from earlier observations. This strategy fully takes into account the multimodality and temporal properties of actions, as well as the differences in pose and gripper states.
Let $\mathcal{O}_{t-k}$ denote the observation at time tkt-k, and let $\mathbf{a}_{t}\mid\mathcal{O}_{t-k}$ represent the model’s prediction for the action at time $t$, based solely on the observation available at time tkt-k. We denote this prediction as
$\mathbf{a}_{t}^{(k)}:=\mathbf{a}_{t}\mid\mathcal{O}_{t-k}.$ (13)
We collect a set of such predictions from the past $N$ time steps
$\mathcal{A}_{t}=\left\{\mathbf{a}_{t}^{(k)}\mid k=0,1,\dots,N-1\right\}.$ (14)
Each $\mathbf{a}_{t}^{(k)}\in\mathbb{R}^{7}$ is decomposed as $\mathbf{a}_{t}^{(k)}=\left[\mathbf{p}_{t}^{(k)},g_{t}^{(k)}\right]$, where $\mathbf{p}_{t}^{(k)}\in\mathbb{R}^{6}$ represents the relative end-effector pose, and $g_{t}^{(k)}\in\mathbb{R}$ denotes the scalar gripper state.
Using the most recent prediction $\mathbf{p}_{t}^{(0)}$ as a reference, we calculate the cosine similarity between each $\mathbf{p}_{t}^{(k)}$ and $\mathbf{p}_{t}^{(0)}$
$\text{sim}^{(k)}=\frac{\mathbf{p}_{t}^{(k)}\cdot\mathbf{p}_{t}^{(0)}}{\|\mathbf{p}_{t}^{(k)}\|\cdot\|\mathbf{p}_{t}^{(0)}\|+\varepsilon},$ (15)
where $\varepsilon=10^{-7}$ is a small constant for numerical stability. To smooth the similarity response and ensure differentiability, we apply a sigmoid transformation with a scaling factor $\alpha>0$
$s^{(k)}=\sigma(\alpha\cdot\text{sim}^{(k)}).$ (16)
This yields similarity weights $s^{(k)}\in(0,1)$ that reflect the semantic alignment between the current and past predictions. We intentionally exclude $g_{t}$ from the cosine similarity computation due to semantic misalignment, magnitude disparity, and behavioral decoupling — as the gripper often functions independently of the end-effector pose (e.g., holding an object steady while the arm moves).
To give higher importance to recent predictions, we apply an exponential decay over time
$d^{(k)}=\exp(-\lambda\cdot k),$ (17)
where $\lambda>0$ is a decay coefficient. The total weight $w^{(k)}$ for each pose is then computed as the product of similarity and temporal decay
$w^{(k)}=\frac{\max(0,\ (1-\beta)\cdot s^{(k)}\cdot d^{(k)})}{\sum_{j=0}^{N-1}\max(0,\ (1-\beta)\cdot s^{(j)}\cdot d^{(j)})},$ (18)
where $\beta\in[0,1)$ is a regularization factor that prevents overconfidence in a small subset of predictions.
The final fused pose is computed as the weighted average
$\hat{\mathbf{p}}_{t}=\sum_{k=0}^{N-1}w^{(k)}\cdot\mathbf{p}_{t}^{(k)}.$ (19)
For the scalar gripper state $g_{t}$, we intentionally do not apply similarity-based weighting, because the cosine similarity measure is computed based on $\mathbf{p}_{t}$, and does not meaningfully reflect the behavior or intent of $g_{t}$. Instead, we use temporal decay alone to perform a weighted average over historical gripper values
$\hat{g}_{t}=\sum_{k=0}^{N-1}\tilde{d}^{(k)}\cdot g_{t}^{(k)},\quad\text{where }\tilde{d}^{(k)}=\frac{d^{(k)}}{\sum_{j=0}^{N-1}d^{(j)}}.$ (20)
The final output action $\hat{\mathbf{a}}_{t}$ is a concatenation of the fused pose and gripper state $\hat{\mathbf{a}}_{t}=[\hat{\mathbf{p}}_{t},\hat{g}_{t}]\in\mathbb{R}^{7}$.
IV Experiment
TABLE III: Comparison results on SIMPLER with WidowX
Method
Fine-tuned
(in-domain)
Put spoon on towel
Put carrot on plate
Stack green block
on yellow block
Put eggplant
on yellow basket
Avg. Grasp Avg. Success
Grasp Success Grasp Success Grasp Success Grasp Success
RT-1-X [7] 16.7 0.0 20.8 4.2 8.3 0.0 0.0 0.0 11.5 1.1
Octo-Base [8] 34.7 12.5 52.8 8.3 31.9 8.3 66.7 43.1 46.5 16.0
Octo-Small [8] 77.8 47.2 27.8 9.7 40.3 4.2 87.5 56.9 58.4 29.5
OpenVLA [9] 4.1 0.0 33.3 0.0 12.5 0.0 8.3 4.1 14.6 1.0
RoboVLMs [12] 70.8 45.8 33.3 20.8 54.2 4.2 91.7 79.2 62.5 37.5
CogACT [13] 71.7 50.8 15.0 67.5 51.8
SpatialVLA [10]
25.0 20.8 41.7 20.8 58.3 25.0 79.2 70.8 51.2 34.4
SpatialVLA [10]
20.8 16.7 29.2 25.0 62.5 29.2 100.0 100.0 53.1 42.7
ThinkAct [27] 58.3 37.5 8.7 70.8 43.8
FPC-VLA 79.2 79.2 58.3 58.3 70.8 45.8 79.2 75.0 71.9 64.6
TABLE IV: Comparison results on SIMPLER with Google robot
Visual Matching Variant Aggregation
Method
Fine-tuned
(in-domain)
Pick Coke
Can (300)
Move
Near (240)
Open/Close
Drawer (216)
Open
Top Drawer
and Place
Apple (108)
Average
Pick Coke
Can (825)
Move
Near (600)
Open/Close
Drawer (378)
Open
Top Drawer
and Place
Apple (189)
Average
RT-1 [5] 85.7 44.2 73.0 6.5 52.4 89.8 50.0 32.3 2.6 43.7
RT-1-X [7] 56.7 31.7 59.7 21.3 42.4 49.0 32.3 29.4 10.1 30.2
RT-2-X [7] 78.7 77.9 25.0 3.7 46.3 82.3 79.2 35.3 20.6 54.4
Octo-Base [8] 17.0 4.2 22.7 0.0 11.0 0.6 3.1 1.1 0.0 1.2
OpenVLA [9] 18.0 56.3 63.0 0.0 34.3 60.8 67.7 28.8 0.0 39.3
RoboVLMs [12] 77.3 61.7 43.5 24.1 41.9
CogACT [13] 91.3 80.5 71.8 50.9 74.8 89.6 80.8 28.3 46.6 61.3
SpatialVLA [10]
81.0 69.6 59.3 89.5 72.7 41.8
SpatialVLA [10]
86.0 77.9 57.4 88.0 72.7 41.8
ThinkAct [27] 92.0 72.4 50.0 84.0 63.8 47.6
FPC-VLA 95.3 93.8 76.4 46.3 78.0 91.3 86.7 30.7 54.5 65.8
TABLE V: Comparison results on LIBERO with Franka
Method Goal Spatial Object Long Average
Diffusion [28] 68.3 78.3 82.5 50.5 72.4
Octo-Base [8] 84.6 78.9 85.7 51.1 75.1
OpenVLA [9] 79.2 84.7 88.4 53.7 76.5
TravceVLA [11] 75.1 84.6 85.2 54.1 74.8
SpatialVLA [10] 78.6 88.2 89.9 55.5 78.1
ThinkAct [27] 87.1 88.3 91.4 70.9 84.4
Cot-VLA [29] 87.6 87.5 91.6 69.0 81.1
FPC-VLA 86.2 87.0 92.0 82.2 86.9
IV-A Implementation Details
The VLM model of FPC-VLA adopts Prismatic VLMs [30]. The VLM model of the supervisor uses Qwen2.5-vl 7B [31]. During training, we use 8 NVIDIA H100 GPUs. For the VLA model, the batch size is set to 256, the learning rate is $2\times 10^{-5}$, and 15 action steps are predicted per inference. For the supervisor model, training is conducted with bfloat16 precision, a batch size of 64, a learning rate of $10^{-4}$, LoRA rank of 8, LoRA alpha of 32, and max pixels of 1003520. During inference, for WidowX, the supervisor is trained on dataset generated from BridgeV2 [25] (408771 entries), with $\delta_{\text{trans}}^{\text{small}},\delta_{\text{trans}}^{\text{large}},\delta_{\text{rot}}^{\text{small}},\delta_{\text{rot}}^{\text{large}}$ set to 0.001, 0.01, 0.001, and 0.01 respectively, and the size of action window $N=7$. For Google Robot, dataset is generated from [5] (728760 entries), with $\delta_{\text{trans}}^{\text{small}},\delta_{\text{trans}}^{\text{large}},\delta_{\text{rot}}^{\text{small}},\delta_{\text{rot}}^{\text{large}}$ set to 0.01, 0.1, 0.01, and 0.1 respectively, and $N=2$. For Franka, dataset is generated from [26] (15604 entries), with $\delta_{\text{trans}}^{\text{small}},\delta_{\text{trans}}^{\text{large}},\delta_{\text{rot}}^{\text{small}},\delta_{\text{rot}}^{\text{large}}$ set to 0.01, 0.1, 0.001, and 0.01 respectively. $N=5$ for the goal setting, $N=7$ for the spatial setting, and $N=11$ for others. In action fusion module, $\alpha=0.1$, $\lambda=\beta=0.01$. Supervisor is called no more than 3 times in each task.
FPC-VLA achieves inference time of 0.176s for non-keyframes and 1.766s for keyframes. As the supervisor is called at most three times during the hundreds of inferences per task, its benefit becomes more evident in long-horizon tasks. For example, in the Open Top Drawer and Place Apple task, the total inference time of FPC-VLA is 40.5 s, which is only slightly higher than the 35.2 s when the supervisor is inactive.
IV-B Evaluation on Simulation Benchmarks
To evaluate the FPC-VLA model, we conduct experiments on two simulation platforms—SIMPLER [32] and LIBERO [26] —as well as on three types of robotic arms: WidowX, Google Robot, and Franka. On SIMPLER, we primarily highlight FPC-VLA’s zero-shot learning capability, using only pre-training on the OXE dataset [7] without domain-specific fine-tuning. On LIBERO, we focus on comparing the performance with other methods after fine-tuning. All evaluation tasks follow [11, 12, 13, 27].
Evaluation on SIMPLER  SIMPLER [32] is an open-source suite of simulated environments to evaluate policies. Visual Matching reduces the visual disparity by overlaying real-world backgrounds and adjusting object textures to align with real environments. Variant Aggregation introduces multiple randomized variations in the simulation’s visual elements. In SIMPLER, we test FPC-VLA on two commonly used robotic platforms: WidowX and Google Robot. Visual results can be found in Fig. 3.
WidowX is tested for grasping, stacking blocks, and placing items into containers. Each task is repeated 24 times under varying object poses. Table III compares various methods with FPC-VLA, reporting performance in success rates (%). RT-1-X [7] and OpenVLA [9] perform poorly across most tasks, with task success rates near 0%, indicating difficulty in generalization due to domain gaps. Notably, FPC-VLA even outperforms the fine-tuned SpatialVLA [10] under the zero-shot setting, demonstrating its effectiveness in both grasping and task completion across varied manipulation scenarios.
Google Robot evaluation includes pick-and-place operations and drawer opening/closing tasks. Each task is repeated under varying object poses and distractors. The number of attempts for each task is indicated in parentheses in Table IV. As shown in Table IV, in Visual Matching category, FPC-VLA achieves an average score of 78.0%, outperforming all baselines, despite in zero-shot setting. For Variant Aggregation tasks, which evaluates the ability to generalize across variants, FPC-VLA again ranks highest. It is worth noting that RT-1 [5] is trained solely on Google Fractal dataset [5], but FPC-VLA significantly outperforms it without any fine-tuning on that dataset. Even when faced with the most complex task (“Open Top Drawer and Place Apple” in Variant Aggregation setting), FPC-VLA still performs remarkably well, surpassing the results of other methods.
Evaluation on LIBERO   LIBERO [26] simulation platform features four distinct task suites: LIBERO-Goal concentrates on achieving different task goals with fixed objects and spatial arrangements. LIBERO-Spatial focuses on adapting to varying spatial configurations of objects while maintaining object consistency. LIBERO-Object involves interacting with a diverse set of objects. LIBERO-Long contains long-horizon tasks that involve diverse objects, layouts, and goals.
Figure 4: Real world experiments results of FPC-VLA on Xiaomi Robot and ALOHA. Supervisor’s failure correction process is demonstrated taking the task “Stack orange block on green block” as an example.
TABLE VI: Results of ablation study
Module Setting
Put spoon on towel
Put carrot on plate
Stack green block
on yellow block
Put eggplant
on yellow basket
Avg. Grasp Avg. Success
Grasp Success Grasp Success Grasp Success Grasp Success
Supervisor w/o Supervisor 83.3 75.0 62.5 50.0 50.0 37.5 79.2 70.8 50.0 58.3
w/o correction 75.0 75.0 58.3 45.8 58.3 45.8 83.3 66.7 68.7 58.3
w/o example samples
75.0 70.8 45.8 41.7 62.5 50.0 83.3 83.3 66.7 61.5
Action Fusion
Module
use latest prediction
29.2 12.5 41.7 8.3 41.7 20.8 16.7 0 32.3 10.4
directly average 79.2 79.2 58.3 54.2 62.5 50.0 83.3 66.7 70.8 62.5
w/o time decay 79.2 70.8 50.0 37.5 58.3 33.3 79.2 66.7 66.7 52.1
motion disturbance + w/o Supervisor 62.5 54.2 50.0 41.7 50.0 20.8 37.5 25.0 50.0 35.4
motion disturbance + FPC-VLA 75.0 62.5 62.5 54.2 62.5 37.5 62.5 50.0 65.6 51.1
FPC-VLA 79.2 79.2 58.3 58.3 70.8 45.8 79.2 75.0 71.9 64.6
Following the data modification and augmentation procedures from [9], we regenerate higher-resolution images (256×256px) , remove no-op actions, rotate third-person images by 180°, filter out failed demonstrations, and use only third-person camera views for fair comparison. Each task suite is evaluated over 500 trials. As shown in Table V, FPC-VLA achieves the highest average performance of 86.9%, significantly outperforming all other baselines. Notably, on the most challenging long-horizon tasks, it scores 82.2%, well above the second-best 70.9% by ThinkAct [27], indicating that failure prediction and correction can help avoid deviation in long-term tasks.
IV-C Real-World Experiments
To assess FPC-VLA’s real-world applicability, as shown in Fig. 4, we evaluate five challenging tasks on the Xiaomi Robot and ALOHA. Xiaomi Robot is a wheeled dual-arm system with 7-DoFs per arm, considering only the gripper’s open/close DoF. ALOHA has dual 6-DoF arms with parallel grippers. For each task, FPC-VLA is fine-tuned with 1000 in-domain demonstrations, while 1000 failure–correction pairs are used to fine-tune the supervisor. Inference uses FP16 on a single NVIDIA RTX 3090, with 32 trials per task. Results show that FPC-VLA achieves robust performance on long-horizon, cluttered tasks, with an average success rate of 86.3%. Notably, we include a dual-arm handover task (Put cup on the left of table), which demonstrates FPC-VLA’s versatility across multi-configuration robots.
IV-D Evaluation of Failure Prediction and Correction
To evaluate the accuracy of our supervisor’s failure prediction and correction, we compare it with several advanced VLMs [33, 34, 35, 31]. Our supervisor is fine-tuned using the failure correction dataset generated as described in Section III-C, while the other VLMs use publicly available pretrained weights. As shown in Fig. 5, our supervisor outperforms all other methods on failure correction datasets derived from three different robotics datasets. Notably, our approach surpasses the base model Qwen2.5-VL-7B using only fine-tuning on generated data.
Figure 5: Evaluation of failure prediction and correction accuracy between our supervisor and other advanced VLMs.
IV-E Ablation Study
Table VI presents the ablation study on FPC-VLA, focusing on the Supervisor and Action Fusion Module. “w/o” indicates “without”. Results are measured on four WidowX tasks. Removing the Supervisor greatly degrades performance, as it corrects deviations from the expert trajectory. Removing corrections during data generation and inference also lowers success, indicating that VLM’s correction suggestions aid action optimization. Providing in-context examples further improves success by clarifying task requirements and output format. For the Action Fusion Module, similarity-guided fusion outperforms both no fusion and simple averaging by better leveraging multi-modal actions, while the time-decay strategy boosts success by emphasizing recent predictions.
To further evaluate the Supervisor’s effect on robustness, we perform an action perturbation experiment with and without the Supervisor. The end-effector pose is randomly disturbed within the range of [0.001, 0.01]. Results show that although perturbations reduce success rates in both settings, the drop is 39.3% without the Supervisor but only 20.9% with FPC-VLA.
V Conclusion
In this paper, we propose FPC-VLA, a VLA Framework with a supervisor for failure prediction and correction. By integrating a VLM-based supervisor and a similarity-guided action fusion module, FPC-VLA not only refines robot actions before execution but also proactively mitigates potential failures. Extensive experiments across simulated environments and real-world robotic platforms demonstrate that FPC-VLA achieves superior performance over existing methods in both zero-shot and fine-tuned settings, highlighting its robustness, generalization ability, and potential for safe and reliable deployment in open-ended robotic applications.